A fair spinner has 9 equal sections: 2 red, 3 blue and 4 green.

It is spun twice.

What is the probability of getting not green then green?


1
Expert's answer
2022-02-27T01:53:02-0500
"P(G^CG)=\dfrac{2+3}{9}(\dfrac{4}{9})=\dfrac{20}{81}"